OpenStudy (anonymous):

give the most general solutions to equations 2sinxcosx-sin(2x)cos(2x)=0 A. simplify the first expression using double angle identity for sine. B. factor the left side of equatioin C. solve the factored equation

OpenStudy (anonymous):

i dont know how to simplify it, factor ir nor solve

OpenStudy (lalaly):

2sinxcosx = sin(2x) 2sinxcosx-sin(2x)cos(2x)=0 sin(2x)-sin(2x)cos(2x)=0 now factor out sin(2x) sin(2x)[1 - cos(2x)] = 0 sin(2x) = 0..since sin(0) = 0.. 2x = 0 x = 0 1- cos(2x) = 0 1 = cos(2x)...since cos(0) =1....2x = 0 x = 0

OpenStudy (mertsj):

\[2\sin x \cos x-2\sin x \cos x(1-2\sin ^2x)=0\] \[2\sin x \cos x[1-(1-2\sin ^2x)]=0\] \[2\sin x \cos x(2\sin^2x)=0\] \[2\sin x=0, \cos x=0,,2\sin ^2x=0\]

OpenStudy (mertsj):

Solve each equation.

OpenStudy (anonymous):

how do you solve for the second equation you have listed?